Washing soda crystals may be used to bleach linen. When crystallized from water, washing soda (sodium carbonate, Na2CO3) forms crystals of a hydrated ionic compound. When 5.00g of washing soda crystals were dried in a desiccator, 1.85g of sodium carbonate remained. Calculate the empirical formula of the hydrated compound.
1
Expert's answer
2020-04-22T09:17:39-0400

Solution.

"Na2CO3 \times nH2O = Na2CO3 + nH2O"

"\frac{5.00}{105.99 + 18.02\times n} = \frac{1.85}{105.99}"

"0.3144 \times n = 3.1505"

n = 10

Answer:

Na2CO3*10H2O